The combustion of octane proceeds according to the following reaction.                 2C8H18(l)+25O2(g)→16CO2(g)+18H2O(l)If 474 moles of octane combust, what volume of carbon dioxide is produced at 19.0 degrees Celsius and 0.995 atm?


1
Expert's answer
2021-06-22T04:42:42-0400

According to the reaction:

n(CO2) = "\frac{16}{2}" n(C8H18) "= 8 \times 474 = 3792 \;mol"

Ideal Gas Law:

"pV=nRT \\

V= \frac{nRT}{p} \\

p=0.995 \;atm \\

T= 19.0 + 273.15 = 292.15 \;K"

R = 0.08206 L×atm/mol×K

"V= \frac{3792 \times 0.08206 \times 292.15}{0.995} = 91365.58 \;L"

Answer: 91365.58 L